OpenStudy (anonymous):

find the zeroes of the function. h(x)=x^2+23x+60

OpenStudy (anonymous):

x^2+23x+60 Delta=23^2-4*1*60 Delta=529-240 Delta= 289 x=(-b+-Sqr(Delta))/2*a x1=(-23+17)/2 x1=-3 x2=(-23-17)/2 x2=-20 x1=-3 or x2= -20